# The multigraded BGG correspondence in Macaulay2

### v1

@@ 1. Introduction
The BGG correspondence admits a generalization involving multigraded polynomial rings, i.e. polynomial rings graded by $\mathbb{Z}^{t}$ for some $t\geq 1$. Just as the classical BGG correspondence plays an important role in projective geometry, the multigraded BGG correspondence plays an analogous role in toric geometry; see e.g. [baranovsky, BETate, BEstrands, BEsyzygies, BEpositivity, BS]. The goal of this article is to describe the Macaulay2 [M2] package MultigradedBGG, which builds on the BGG package by implementing the multigraded BGG correspondence.
@@ Acknowledgments
We are grateful to Daniel Erman and Keller VandeBogert for generously contributing many key insights to the code in MultigradedBGG and to Greg Smith for many helpful comments during the writing of this package. We thank the Institute for Mathematics and its Applications at the University of Minnesota for hosting the “Macaulay2 Workshop & Mini-School” (funded by NSF grant DMS-2302476) on June 3 - 9, 2023, during which most of this package was developed.
@@ 2. Computing with differential modules
A straightforward calculation shows that the differential module $(D,\partial_{D})$ from Example 2.2 is not isomorphic to a free flag. However, there is a quasi-isomorphism $(G,\partial_{G})\xrightarrow{\simeq}(D,\partial_{D})$ [BE1, Example 5.8].
@@
Since we may replace $D$ with $D(n)$, we may assume $n=0$. We proceed via the following algorithm, which is very similar to Algorithm 2.7:
@@
When a differential module $D$ admits a minimal free flag resolution, one may also obtain it by first resolving $D$ via Algorithm 2.7 and then minimizing, but this is less efficient than Algorithm 2.11. This algorithm is implemented in MultigradedBGG as resMinFlag.
@@ 4. Computing strongly linear strands of multigraded free resolutions
As in Examples 3.1 and 3.2, let $S=k[x_{0},\dots,x_{3}]$ be the $\mathbb{Z}^{2}$-graded Cox ring of the Hirzebruch surface of type 3. The minimal free resolution of $M=S/(x_{1},x_{2}^{2})$, is, of course, the Koszul complex on $x_{1}$ and $x_{2}^{2}$. The strongly linear strand of this resolution is therefore the Koszul complex on $x_{1}$. Let us verify this with MultigradedBGG:

### v2

@@ 1. Introduction
The BGG correspondence admits a generalization involving multigraded polynomial rings, i.e. polynomial rings graded by $\mathbb{Z}^{t}$ for some $t\geq 1$. Just as the classical BGG correspondence plays an important role in projective geometry, the multigraded BGG correspondence plays an analogous role in toric geometry; see e.g. [baranovsky, BETate, BEstrands, BEsyzygies, BEpositivity, BS, EES]. The goal of this article is to describe the Macaulay2 [M2] package MultigradedBGG, which builds on the BGG package by implementing the multigraded BGG correspondence.
@@ Acknowledgments
We are grateful to Daniel Erman and Keller VandeBogert for generously contributing many key insights to the code in MultigradedBGG and to Greg Smith for many helpful comments during the writing of this package. We thank the Institute for Mathematics and its Applications at the University of Minnesota for hosting the “Macaulay2 Workshop & Mini-School” (funded by NSF grant DMS-2302476) on June 3 - 9, 2023, during which most of this package was developed. Finally, we thank the referees for many helpful suggestions.
@@ 2. Computing with differential modules
The differential module $(D,\partial_{D})$ from Example 2.2 is not isomorphic to a free flag [ABI, Example 5.6]. But, there is a quasi-isomorphism $(G,\partial_{G})\xrightarrow{\simeq}(D,\partial_{D})$ [BE1, Example 5.8].
@@
Applying the function differentialModule to an object of type Matrix yields an object of type DifferentialModule, which is a Complex with three properties: it is only nonzero in homological degrees $-1$, 0, and 1; its two differentials are identical; and the differentials compose to zero.
@@
Since we may replace $D$ with $D(n)$, we may assume $n=0$. We proceed via the following algorithm:
@@
Algorithm 2.11 is implemented in MultigradedBGG as resMinFlag. When a differential module $D$ admits a minimal free flag resolution, one may also obtain it by first resolving $D$ via Algorithm 2.7 and then minimizing, but this is less efficient than Algorithm 2.11.
@@ 3. The multigraded BGG functors
Example 3.3.
Suppose $S=k[x_{0},\dots,x_{n}]$ is $\mathbb{Z}$-graded, where $\deg(x_{i})=d_{i}\geq 1$. In other words, $S$ is the Cox ring of the weighted projective space $X=\mathbb{P}(d_{0},\dots,d_{n})$. Given a finitely generated $S$-module $M$, it follows from [BETate, Theorem 3.7] that the cohomology of the twists of the sheaf $\widetilde{M}$ on $X$ may be computed by taking a minimal free resolution of $\mathbf{R}(M)$. In more detail: let $\varepsilon\colon F\xrightarrow{\simeq}\mathbf{R}(M)$ be the minimal free resolution of $\mathbf{R}(M)$, and let $T(\widetilde{M})$ be the mapping cone of $\varepsilon$. The differential $E$-module $T(M)$ is called the Tate resolution of the sheaf $\widetilde{M}$. By [BETate, Theorem 3.7], we have:
(3.1) $H^{j}(X,\widetilde{M}(i))\cong\operatorname{Hom}_{E}(k,T(\widetilde{M}))_{(i,-j)}.$
As a simple example, suppose $X=\mathbb{P}(1,1,2)$ and $M=S/(x_{0},x_{1})$. We have
(3.2) $\dim_{k}H^{j}(X,\widetilde{M}(i))=\begin{cases}1,&j=0\text{ and }i\text{ even};\\ 0,&\text{else}.\end{cases}$
Let us check this via the isomorphism (3.1). We have $\mathbf{R}(M)=\bigoplus_{i\geq 0}\omega_{E}(-2i,0)$, with differential mapping $\omega_{E}(-2i,0)$ to $\omega_{E}(-2i-2,0)$ via multiplication by $e_{2}$. The minimal free resolution of $\mathbf{R}(M)$ is easily seen to be $F=\bigoplus_{i\geq 1}\omega_{E}(2i,1)$, with differential that sends $\omega_{E}(2,1)$ to zero and $\omega_{E}(2i,1)$ to $\omega_{E}(2i-2,1)$ via multiplication by $e_{2}$ for $i\geq 2$. As an $E$-module, we have $T(\widetilde{M})=F(0,-1)\oplus\mathbf{R}(M)=\bigoplus_{i\in\mathbb{Z}}\omega_{E}(2i,0).$ The calculation (3.2) thus follows from the isomorphism (3.1).
One may use MultigradedBGG to check all of this, but there are subtleties to be aware of. We first compute $\mathbf{R}(M)$ (in a finite window):
i1 : X = weightedProjectiveSpace {1,1,2};
i2 : S = ring X;
i3 : M = coker matrix{{x_0, x_1}};
i4 : D = toricRR(M, for i from 0 to 4 list i);
i5 : D.dd_0
o5 = {4, 3} | 0   0   0 |
     {6, 3} | e_2 0   0 |
     {8, 3} | 0   e_2 0 |
Let us now resolve $\mathbf{R}(M)$. While the function resDM does not always yield a minimal free resolution, it does give a portion of the minimal free resolution of $\mathbf{R}(M)$ in this case (the problem of non-minimality can be solved by slightly modifying Algorithm 2.11, but we will not pursue this here). Since we have truncated $\mathbf{R}(M)$, we have created additional homology that manifests itself as extraneous summands in the free resolution.
i6 : F = resDM(D, 3);
i7 : d = F.dd_0
o7 = {8, 3}  | 0 0 e_2 0   0   0   |
     {2, 2}  | 0 0 0   e_2 0   0   |
     {6, 3}  | 0 0 0   0   e_2 0   |
     {0, 2}  | 0 0 0   0   0   e_2 |
     {4, 3}  | 0 0 0   0   0   0   |
     {-2, 2} | 0 0 0   0   0   0   |
The submatrix corresponding to the degree $(2,2),(0,2)$, and $(-2,2)$ generators gives the summand we want; the rest corresponds to homology of our truncation of $\mathbf{R}(M)$ that we wish to ignore.
i9 : submatrix’(d,{0,2,4},{0,2,4})
o9 = {2, 2}  | 0 e_2 0   |
     {0, 2}  | 0 0   e_2 |
     {-2, 2} | 0 0   0   |
Since these three summands have socle degree $(-2,-1),(-4,-1)$, and $(-6,-1)$, their presence implies that $H^{0}(X,\widetilde{M}(j))=1$ for $j=-2,-4,-6$. To compute the cohomology of additional negative twists, one needs to resolve a larger portion of $\mathbf{R}(M)$.
In general, to compute the cohomology of a sheaf $\widetilde{M}$ on weighted projective space using the methods in MultigradedBGG, there are delicate matters to address. For instance, what is the most efficient window of $\mathbf{R}(M)$ to use? How much of the (typically infinite) free resolution of $\mathbf{R}(M)$ must one compute in order to be sure one has included all of the needed summands? How does one ensure that the extraneous summands resulting from the truncation of $\mathbf{R}(M)$ do not affect the calculation? All of these matters will be addressed in future work of the second author with Daniel Erman.
@@ 4. Computing strongly linear strands of multigraded free resolutions
As in Examples 3.1 and 3.2, let $S=k[x_{0},\dots,x_{3}]$ be the $\mathbb{Z}^{2}$-graded Cox ring of the Hirzebruch surface of type 3. The minimal free resolution of $M=S/(x_{0},x_{1}^{2})$, is, of course, the Koszul complex on $x_{0}$ and $x_{1}^{2}$. The strongly linear strand of this resolution is therefore the Koszul complex on $x_{0}$. Let us verify this with MultigradedBGG: